OpenStudy (anonymous):

Please help! what is the trigonometric form of -4i?

OpenStudy (anonymous):

ln[cos(4)-isin(4) that is my guess

OpenStudy (anonymous):

These are my options 4(cos 180° + i sin 180°) 4(cos 270° + i sin 270°) 4(cos 90° + i sin 90°) 4(cos 0° + i sin 0°)

OpenStudy (anonymous):

I used \[e ^{i \theta}=\cos(\theta)+isin(\theta)\]

OpenStudy (anonymous):

which one would that be?

OpenStudy (anonymous):

b looks like the only one that is right

OpenStudy (anonymous):

Thank you

OpenStudy (anonymous):

You need to know what values on the unit circle each degree equals for the sin and cos components

OpenStudy (anonymous):

will you help with another?

OpenStudy (anonymous):

this one is harder

OpenStudy (anonymous):

sure make another thread and call me over

OpenStudy (anonymous):

alright 1 second